OpenStudy (anonymous):

The probabilities of the events A and B are 0.20 and 0.30, respectively. The probability that both A and B occur is 0.15. What is the probability of either A or B occurring.

OpenStudy (anonymous):

P(A or B) = P(A) + P(B), assuming A and B are mutually exclusive. Hence, P(A or B) = .20 + .30 = .50 or 1/2.

OpenStudy (anonymous):

P(A or B) = P(A ) + P (B) -P(A and B) hence P(A+B) =.2+.3-.15 = .35

OpenStudy (anonymous):

this is because you would be counting twice the intersection of the two events